# CALT-TH-2024-013

### v2

@@ Abstract
Motivated by string-theoretic swampland conjectures, the existence of a dark fifth dimension, whose size is roughly 1 – 10 microns, has been proposed. A great deal of supporting evidence has been presented, and definitive experimental tests are likely to be carried out. The basic idea is that the four-dimensional spacetime that we observe is confined to a brane that is localized in the dark dimension. This short note points out that there are two distinct ways to realize such a scenario in string theory/M-theory. In the one considered previously the dark dimension is topologically a circle and our observable 4d spacetime is confined to a brane that is localized in a GUT-scale region of the circle. An alternative possibility is that the dark dimension is a line interval with branes at each end. The latter option would imply the existence of a parallel 4d spacetime microns away from us!
Motivated by string-theoretic swampland conjectures [1], it has recently been proposed that there is a mesoscopic fifth dimension whose size is roughly 1 – 10 microns [2][6]. One aspect of this is the proposal that a tower of light states in our universe is related to the smallness of the cosmological constant $\Lambda$ [7]. The new idea is that this tower corresponds to Kaluza–Klein (KK) modes of a graviton in a dark dimension, and that they constitute dark matter. The size of this dimension is of order $\Lambda^{-1/4}$. In this way dark energy and dark matter are connected. This dimension is called dark, because it only supports gravitational forces. The basic idea is that the four-dimensional spacetime that we observe lives on a brane that is localized at a point (more precisely, a GUT-scale region) in the dark dimension (DD). Any additional compact dimensions must be very much smaller than the DD. Thus, for many purposes, we may assume that the brane is well approximated by a standard-model 3-brane (SM3-brane) even though it would be desirable to understand exactly how such a brane arises from a more fundamental string theory/M-theory starting point. For reasons explained in [2], the existence of more than one mesoscopic dimension is excluded. As discussed in [3], astrophysical constraints limit the allowed size of the DD to 1 – 10 microns, and future large-scale structure surveys could support or rule out its existence.
@@
There have been previous theoretical proposals for a relatively large extra dimension, such as [15] and [16]. They differ in important respects from the current proposal, which is motivated by swampland considerations as well as various astrophysical arguments. In particular, the current proposal shows promise for giving a deeper quantitative understanding of dark matter and dark energy without invoking the anthropic principle.
@@
An F-theory approach to realizing the DD proposal is described in [6]. In that realization the DD is topologically a circle, and the SM3-brane is localized in a GUT-scale region of the circle. That paper argues that this picture gives a tower of states, associated to a 7-brane, at $10^{9}$$10^{10}$ GeV. This scale, called the species scale, corresponds to the 5d Planck scale. There should also be a GUT scale a few orders of magnitude higher. The SM3-brane approximation should be sufficient for phenomenology in the foreseeable future, though a deeper understanding of how the SM3-brane arises would be desirable.
@@
Among the many exciting discoveries made in the mid-1990s was the realization that the $E_{8}\times E_{8}$ heterotic string theory and the type IIA superstring theory actually have an eleventh dimension that is not visible in perturbation theory [22][23]. In the type IIA case there is a circular eleventh dimension ($S^{1}$) whose size is proportional to the string coupling constant. In the limit of infinite coupling one obtains 11-dimensional M-theory. In the case of the $E_{8}\times E_{8}$ theory there is a orbifold eleventh dimension ($S^{1}/Z_{2}$), whose size is proportional to the string coupling constant. This dimension can also be viewed as a line interval with two ends. Each of the ends is attached to an end-of-the-world 9-brane (an EW9-brane). One $E_{8}$ gauge field is associated to each of the two EW9-branes. This is required by the anomaly analysis [22]. The 11-dimensional bulk, in between the EW9-branes, is described by 11-dimensional M-theory, which is the UV completion of 11d supergravity [24].
@@
The eleventh dimension of the $E_{8}\times E_{8}$ theory is a DD. Most studies of Calabi–Yau compactification of this theory have ignored the eleventh dimension, which may be a reasonable thing to do for many purposes. However, it is an important fact that this class of string theory solutions necessarily implies the existence of a DD. The DD has been taken into account in [25][28]. As explained in [25], one consequence of going to strong coupling is to reduce the lower bound on the 4d Newton constant $G_{N}$. This is important, since $G_{N}$ is too large in the weakly coupled $E_{8}\times E_{8}$ theory. So the requirement that the DD in this theory must be large has been understood for a long time. What is new now is the suggestion, based on swampland considerations, that the size of this dimension should be several microns.
@@
Suppose we consider implementing the DD proposal using the $E_{8}\times E_{8}$ theory. Then an essential difference from the F-theory example is that it requires the existence of two EW-branes. After a suitable compactification of six dimensions they can be approximated by 3-branes, and one of them can be identified as the SM3-brane that describes our 4d reality at least up to energies of a TeV and possibly much further. In this setup there is a second parallel 3-brane, just microns away! Should we expect the physics of the second EW-brane to be the same as or different from “our” EW-brane? Much of the Calabi–Yau phenomenology has found it useful to implement supersymmetry breaking in the invisible $E_{8}$, by gaugino condensation for example, and for this breaking to be transferred to the other observable $E_{8}$ factor gravitationally. This implies that the physics of the two EW-branes is different. Whether or not that is the case, the possibility of a whole other parallel 4d spacetime just microns away from us is challenging to visualise and seems quite incredible! This is not a prediction of the DD proposal – just one of the two options.
@@
For strong coupling the $E_{8}\times E_{8}$ heterotic string theory is not a string theory. (A similar remark can be made for the type IIA superstring theory.) In M theory the basic branes are an M2-brane and its magnetic dual, which is an M5-brane. A cylindrical M2-brane can be stretched across the dark dimension with its boundaries attached to the EW-branes. In the weak-coupling limit the two EW-branes approach one another, and the cylinder collapses to become the $E_{8}\times E_{8}$ string. However, if the branes are microns apart, it looks nothing like a string – unless it can be approximated by an open string stretched between the two EW-branes. In the $E_{8}\times E_{8}$ theory the coupling constant $g$ is proportional to the radius of the dark dimension, $\rho$. More specifically, $g\sim\rho/l_{s}$, where $l_{s}$ is the string scale, which is very small. If $\rho$ is several microns, this ratio is enormous. The problem is evident from the shape of the stretched M2-brane. If there were an S-dual formulation of the $E_{8}\times E_{8}$ heterotic string theory, whose coupling constant is $\tilde{g}=1/g$, that could be helpful. The $SO(32)$ heterotic string theory has an S-dual formulation, namely the type I superstring theory, which is based on a non-BPS (breakable) string. It seems unlikely that there is an analogous S-dual of the $E_{8}\times E_{8}$ heterotic string theory. One problem is that it seems to be impossible to define a $g\to\infty$ limit while retaining the EW-branes.
@@
Alternatively, it is conceivable that a very sensitive test of gravity at short distances could distinguish between the two geometries. If the circle has radius $\rho=1/\mu$, then summing Yukawa potentials for the KK excitations gives a potential proportional to
@@
The leading large $r$ asymptotics are shown. For small $r$ they are both proportional to $1/r^{2}$, as expected, but differ by a factor of 2.
There is a subtlety here. A massless graviton in 5d has five polarizations, forming an irreducible representation of the SU(2) little group, as does a massive spin two particle in 4d. The only mismatch here concerns the zero modes. In both geometries there are three extra massless modes in 4d with helicities +1, 0, -1 in addition to the desired graviton with helicities +2 and -2. They are unwanted, since it is well established that 4d gravity is mediated by a massless spin 2 graviton.
A plausible resolution of this problem is that these three modes combine to give a massive spin one particle in 4d. The mass of the vector particle must go to zero in the decompactification limit, if one is to recover 11d M theory. Therefore, it ought to be proportional to $\mu$. The formulas given above may need to be modified in order to correctly account for this massive vector. In principle, an observer on the SM3-brane should be able to detect one massive spin one particle in addition to the infinite tower of massive spin two particles by carefully measuring the strength of gravity at short distances.
The original papers [2][6] make a strong case for KK excitations of gravitons in the DD comprising the most important component of dark matter. Of course, it is still possible that there are also other sources of dark matter. In addition to the massive spin-one particle discussed in the preceding paragraph, other sources of dark matter in the DD might include KK excitations of sterile neutrinos or other bulk degrees of freedom. Possible sources of dark matter on the SM3-brane are the familiar ones that have received the most attention by both theorists and experimentalists so far. These include axions and wimps, though the latter have been largely excluded. If there is a second EW-brane, matter that is confined to that brane would be dark matter from our point of view.
@@
In conclusion, the DD scenario is a well-motivated proposal, based on plausible string theory considerations, that should be experimentally testable. Its confirmation would be revolutionary for particle physics and astrophysics. If that happens, an important question would be whether or not there is a parallel brane microns away from our SM3-brane.
@@ Acknowledgments
I am grateful to Cumrun Vafa for encouraging me to release this note and to Temple He for a helpful discussion. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award Number DE-SC0011632.

### v3

@@ Abstract
Motivated by string-theoretic swampland conjectures, the existence of a dark fifth dimension, whose size is roughly 1 – 10 microns, has been proposed. A great deal of supporting evidence has been presented, and definitive experimental tests are likely to be carried out. The basic idea is that the four-dimensional spacetime that we observe is confined to a brane that is localized in the dark dimension. This short note points out that there are two distinct ways to realize such a scenario in string theory/M-theory. In the one considered previously, the dark dimension is topologically a circle and our observable 4d spacetime is confined to a brane that is localized in a GUT-scale region of the circle. An alternative possibility is that the dark dimension is a line interval with “end-of-the-world-branes” at each end. The latter option would imply the existence of a parallel 4d spacetime, with the same gravitational forces as ours, but otherwise its own physical laws, microns away from us!
Motivated by string-theoretic swampland conjectures [1], it has recently been proposed that there is a mesoscopic fifth dimension whose radius $\rho$ is roughly 1 – 10 microns [2][6]. One aspect of this is the proposal that a tower of light states in our universe is related to the smallness of the cosmological constant $\Lambda$ [7]. The new idea is that this tower corresponds to Kaluza–Klein (KK) modes of a graviton in a dark dimension, and that they constitute dark matter. The size of this dimension is of order $\Lambda^{-1/4}$. (We always set $\hbar=c=1$.) The unknown coefficient is responsible for the uncertainty in the prediction for $\rho$. In this way dark energy and dark matter are connected without invoking an anthropic principle. This dimension is called dark, because it only supports gravitational forces. The basic idea is that the four-dimensional spacetime that we observe lives on a brane that is localized at a point (more precisely, a GUT-scale region) in the dark dimension (DD). Any additional compact dimensions must be very much smaller than the DD. Thus, for many purposes, we may assume that the brane is well approximated by a standard-model 3-brane (SM3-brane) even though it would be desirable to understand exactly how such a brane arises from a more fundamental string theory/M-theory starting point. For reasons explained in [2], the existence of more than one mesoscopic dimension is excluded. As discussed in [3], astrophysical constraints limit the allowed size of the DD to 1 – 10 microns, and future large-scale structure surveys could support or rule out its existence.
@@
There have been previous theoretical proposals for a relatively large extra dimension, such as [15] and [16]. They differ in important respects from the current proposal, which is motivated by swampland considerations as well as various astrophysical arguments. In particular, the current proposal shows promise for giving a deeper understanding of why the amounts of dark matter and dark energy are comparable without invoking the anthropic principle.
@@
An F-theory approach to realizing the DD proposal is described in [6]. In that realization the DD is topologically a circle, and the SM3-brane is localized in a GUT-scale region of the circle. Among other things, that paper argues that this picture gives a tower of states, associated to a 7-brane, at $10^{9}$$10^{10}$ GeV. This scale, called the species scale, corresponds to the 5d Planck scale.
@@
Among the many exciting discoveries made in the mid-1990s was the realization that the $E_{8}\times E_{8}$ heterotic string theory and the type IIA superstring theory actually have an eleventh dimension that is not visible in perturbation theory [22][23]. In the type IIA case there is a circular eleventh dimension ($S^{1}$) whose size is proportional to the string coupling constant. In the limit of infinite coupling one obtains 11-dimensional M-theory. In the case of the $E_{8}\times E_{8}$ theory there is an orbifold eleventh dimension ($S^{1}/Z_{2}$), whose size also is proportional to the string coupling constant. This dimension can also be viewed as a line interval with two ends. Each of the ends is attached to an “end-of-the-world” 9-brane (an EW9-brane). One $E_{8}$ gauge field is associated to each of the two EW9-branes. This is required by the anomaly analysis [22]. The 11-dimensional bulk, in between the EW9-branes, is described by 11-dimensional M-theory, which is the UV completion of 11d supergravity [24].
@@
The eleventh dimension of the $E_{8}\times E_{8}$ theory is a DD. Most studies of Calabi–Yau compactification of this theory have ignored the eleventh dimension, which may be a reasonable thing to do for many purposes. However, it is an important fact that this class of string theory solutions necessarily implies the existence of a DD. Some of the implications of the DD have been discussed in [25][30]. As explained in [25], one consequence of increasing the string coupling constant (and hence the size of the DD) is to reduce the lower bound on the 4d Newton constant $G_{N}$. This is important, since $G_{N}$ is necessarily too large in the weakly coupled $E_{8}\times E_{8}$ theory. So the requirement that the DD in this theory must be quite large has been understood for a long time. What is new now is the suggestion, based on swampland considerations, that the size of this dimension should be several microns. By the standards of particle physics, this is very large indeed.
@@
Suppose we consider implementing the DD proposal using the $E_{8}\times E_{8}$ theory. Then an essential difference from the F-theory example is that it requires the existence of two EW-branes. After a suitable compactification of six dimensions they can be approximated by 3-branes, and one of them can be identified as the standard model 3-brane (SM3-brane) that describes our 4d reality at least up to energies of a TeV and possibly much further. In this setup there is a second parallel 3-brane, just microns away! Should we expect the physics of the second EW-brane to be the same as or different from “our” EW-brane? The Calabi–Yau phenomenology works best when the two $E_{8}$’s are treated differently. This implies that the physics of the two EW-branes is different, so that seems likely to be the case. In any case, the possibility of a parallel 4d spacetime just microns away from us is challenging to visualise and seems quite incredible! However, as the F-theory example illustrates, this is not an unavoidable prediction of the DD proposal – just one of the two options.
@@
For strong coupling the $E_{8}\times E_{8}$ heterotic theory is not a string theory. (A similar remark can be made for the type IIA theory.) In M theory the basic branes are an M2-brane and its magnetic dual, which is an M5-brane. A cylindrical M2-brane can be stretched across the dark dimension with its boundaries attached to the EW-branes. In the weak-coupling limit the two EW-branes approach one another, and the cylinder collapses to become the $E_{8}\times E_{8}$ string. However, if the branes are microns apart, it looks nothing like a string – unless it can be approximated by an open string stretched between the two EW-branes, but I am unaware of any such proposal. In the $E_{8}\times E_{8}$ theory the coupling constant $g$ is proportional to the radius of the dark dimension, $\rho$. More specifically, $g\sim\rho/l_{s}$, where $l_{s}$ is the string scale, which is very small. If $\rho$ is several microns, this ratio is enormous. The problem is evident from the shape of the stretched M2-brane. If there were an S-dual formulation of the $E_{8}\times E_{8}$ heterotic string theory, whose coupling constant is $\tilde{g}=1/g$, that could be helpful. The $SO(32)$ heterotic string theory has an S-dual formulation, namely the type I superstring theory, which is based on a non-BPS (breakable) string. It seems unlikely that there is an analogous S-dual of the $E_{8}\times E_{8}$ heterotic string theory. One problem is that it is not possible to define a $g\to\infty$ limit while retaining both of the EW-branes, and we don’t want $\rho$ to be infinite.
@@
It is conceivable that a very sensitive test of gravity at short distances could distinguish between the two geometries. If the DD is a circle of radius $\rho=1/\mu$, then summing Yukawa potentials for the KK excitations of a 5d graviton gives a potential proportional to
@@
The leading large $r$ asymptotic behavior is shown. For small $r$ it is proportional to $\rho/r^{2}$ in both cases, as expected, though they differ by a factor of 2.
These formulas probably require corrections for the following reason. A massless graviton in 5d has five polarizations, forming an irreducible representation of the SU(2) little group, as does a massive spin two particle in 4d. The only mismatch here concerns the zero modes. In both cases there are three extra massless modes in 4d with helicities $+1$, 00, $-1$ in addition to the desired graviton with helicities $+2$ and $-2$. They are not wanted, since it is well established that 4d gravity is mediated by a massless spin 2 graviton. A plausible resolution of this problem is that these three modes combine to give a massive spin one particle in 4d. The mass of this vector particle must go to zero in the decompactification limit, if one is to recover 11d M theory. Therefore, it ought to be proportional to $\mu$. The formulas above need to be modified in order to correctly account for this massive vector. In principle, the experiments should be able to reveal one massive spin one particle in addition to the usual graviton and the infinite tower of massive spin two particles by measuring the strength of gravity at short distances to high precision.
An analogous problem occurs when M theory has a circular dimension giving strongly coupled type IIA superstring theory. The type IIA massless spectrum in 10d contains a graviton, a vector, and a scalar that descend from the 11d graviton. The presence of the massless vector reflects the $U(1)$ isometry of the eleventh dimension, in contrast to the cases we have discussed in which one or two branes break the symmetry. In the type IIA case there is a known deformation of the theory in which the scalar and vector combine to give a massive vector [31]. This mass arises when one or more 8-branes are added to the type IIA background [32]. Thus, it is clear that the vector must be massive whenever branes break the $U(1)$ symmetry of the dark dimension. This is the case for the F-theory and M-theory examples under consideration here.
The original papers [2][6] make a strong case for KK excitations of gravitons in the DD comprising the most important component of dark matter. Of course, it is still possible that there are also other sources of dark matter. Other sources of dark matter in the DD might include KK excitations of sterile neutrinos or other bulk degrees of freedom. Possible sources of dark matter on the SM3-brane are the familiar ones that have received the most attention by both theorists and experimentalists so far. These include axions and wimps, though the latter have been largely excluded. If there is a second EW-brane, matter that is confined to that brane would also be dark matter.
@@
One may wonder whether the DD proposal has any bearing on the $\Lambda$CDM model of cosmology. In particular, the time evolution of various quantities is important. Since $G_{N}\sim G_{5}/\rho$, where $G_{5}$ is the 5d Newton constant, and $\Lambda\sim\rho^{-4}$, it follows that
$\frac{\dot{G_{N}}}{G_{N}}=\frac{\dot{G_{5}}}{G_{5}}-\frac{\dot{\rho}}{\rho}{\quad{\rm and}\quad}\frac{\dot{\Lambda}}{\Lambda}=-4\frac{\dot{\rho}}{\rho}.$ (3)
Variation of $G_{N}$ is constrained by various observations, such as binary pulsars, and variation of $G_{5}$ seems unlikely in the epoch where the $\Lambda$CDM model is supposed to apply. Therefore significant variation of $\rho$ and $\Lambda$ appears to be unlikely in this approach. Even so, the DD model may help to address the current stress in the $\Lambda$CDM model. An additional element is the fact that the DD model predicts that Kaluza–Klein excitations of the graviton constitute a spectrum of decaying dark matter particles. While there have been studies of of decaying dark matter as a possible mechanism to address the stress, the details of how it occurs in the DD model are different from the way it has been modeled in these studies. There may also be other effects, such as Casimir forces, that are relevant.
In conclusion, while the existence of a dark dimension of the sort discussed here is certainly speculative, it is a well-motivated and experimentally testable proposal. It is based on plausible string theory considerations that should be experimentally testable. Its confirmation would be revolutionary for particle physics and astrophysics. If the existence of a DD is confirmed, an important question would be whether or not there is a parallel 3-brane microns away from the SM3-brane on which we reside.
@@ Acknowledgments
I am grateful to Cumrun Vafa and Temple He for helpful discussions. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award Number DE-SC0011632.